OpenStudy (anonymous):

Question about Solving Systems. Multiple Choice Question! What is the solution of this system? y = 5x + 4 y = 5x – 3 Thanks! a. no solutions b.(0,4) c. infinitely more solutions d. (0,-3)

hartnn (hartnn):

hint : subtract those 2 equations, what u get ?

OpenStudy (anonymous):

How? I am really confused.

hartnn (hartnn):

to subtract 2 equations, you subtract their left side and put it in left side of = sign and subtract their right side and put it in right side of = sign example y= mx+c y= px+q subtracting these equation, i get y-y = mx+c - px- q so., now u try for y = 5x + 4 y = 5x – 3

OpenStudy (anonymous):

Thanks. I got it, I understand what my textbook was saying now.

hartnn (hartnn):

what u got answer as ?